# The Sunburst Arc with JWST: III. An Abundance of Direct Chemical Abundances

### v2

@@ 6 Conclusions
We have measured direct gas-phase chemical abundances in the stacked spectrum of the Lyman-continuum emitting star cluster in the Sunburst Arc. We detect the [Sii], [Oiii], [Siii], and [Oii] lines, allowing measurement of the electron temperature in the low, intermediate, and high ionization zones of the nebular gas. With these multiple temperature measurements, we measure direct abundances of oxygen, nitrogen, sulfur, argon, neon, and iron, following best practices established in nearby galaxies from the CHAOS program. These measurements enable a direct comparison between well-studied nearby HII regions and an object at Cosmic Noon ().
These high-quality abundance measurements enable detailed studies of the origins of abundance patterns in distant galaxies. In the Sunburst LCE, the combination of these abundance measurements and the detection of Wolf-Rayet features (Rivera-Thorsen et al., 2024) allow us to conclude that nitrogen-rich Wolf-Rayet stars are responsible for increasing the nitrogen abundance.
Further high-quality abundance measurements of galaxies and galactic substructures at Cosmic Noon and beyond will continue to drive our understanding of how galaxies built up the chemical elements.

### v3

@@ 6 Conclusions
We have measured direct gas-phase chemical abundances in the stacked spectrum of the Lyman-continuum emitting star cluster in the Sunburst Arc. We detect the [Sii], [Oiii], [Siii], and [Oii] lines, allowing measurement of the electron temperature in the low, intermediate, and high ionization zones of the nebular gas. With these multiple temperature measurements, we measure direct abundances of oxygen, nitrogen, sulfur, argon, neon, and iron, following best practices established in nearby galaxies from the CHAOS program. These measurements enable a direct comparison between well-studied nearby Hii regions and an Hii region at Cosmic Noon ().
These high-quality abundance measurements enable detailed studies of the origins of abundance patterns in distant galaxies and Hii regions. In the Sunburst LCE, the combination of these abundance measurements and the detection of Wolf-Rayet features (Rivera-Thorsen et al., 2024) allow us to conclude that nitrogen-rich Wolf-Rayet stars are responsible for increasing the nitrogen abundance.
Further high-quality abundance measurements of galaxies and galactic substructures at Cosmic Noon and beyond will continue to drive our understanding of how galaxies and their star clusters built up the chemical elements. For example, further exploration of both temperature and abundance variation in individual Hii regions in a larger sample of gravitationally lensed galaxies could constrain globular cluster formation models, recalibrate empirical temperature relations for high- studies, and examine abundance gradients to constrain galaxy formation models. This study could thus be seen as a proof of concept for future work characterizing abundances of lensed Hii regions.